Please show me how to answer this problem.
3x(2x-1)=0

Answer by askmemath(368)

We know that any number multiplied by 0 is always 0
SO for 3x(2x-1)=0 either 3X = 0 or 2X-1 = 0
If 3X = 0 then X = 0 since 3 does not equal 0
If 2x-1 = 0
Adding 1 on both sides
2X = 1
Dividing by 2 throughout
X = 1/2